OpenStudy (anonymous):

What is the graph of 4x – 6y = 24?

OpenStudy (anonymous):

VHS_PA_S2_02_L208_L307_Q8a.gif A. VHS_PA_S2_02_L208_L307_Q8b.gif B. VHS_PA_S2_02_L208_L307_Q8c.gif C. VHS_PA_S2_02_L208_L307_Q8d.gif D.

OpenStudy (anonymous):

Easy

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

lets put this in y=mx+b form

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

4x – 6y = 24 subtract 4x -6y=-4x+24 divide by -6 y=2/3x-4

OpenStudy (anonymous):

The graph looks like this...

OpenStudy (anonymous):

they all have A B C D under them

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

let me compare them

OpenStudy (anonymous):

so...... C right CX

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

THx

OpenStudy (anonymous):

Ur welcome